### v1

@@ Abstract
Triangle counting in a graph is a fundamental problem and has a wide range of applications in various domains. It is crucial in understanding the structural properties of a graph and is often used as a building block for more complex graph analytics. In this paper, we solve the triangle counting problem in an anonymous graph in a distributed setting using mobile agents and subsequently use this as a subroutine to tackle the truss decomposition and triangle centrality problem. The paper employs mobile agents, placed on the nodes of the graph to coordinate among themselves to solve the triangle enumeration problem for the graph. Following the literature, we consider the synchronous systems where each robot executes its tasks concurrently with all others and hence time complexity can be measured as the number of rounds needed to complete the task. The graph is anonymous, i.e., without any node labels or IDs, but the agents are autonomous with distinct IDs and have limited memory. Agents can only communicate with other agents locally i.e., if and only if they are at the same node. The goal is to devise algorithms that minimise both the time required for triangle counting and the memory usage at each agent. We further demonstrate how the triangle count obtained through the mobile agent approach can be leveraged to address the truss decomposition, triangle centrality and local clustering coefficient problems, which involves finding maximal sub-graphs with strong interconnections. Truss decomposition helps in identifying maximal, highly interconnected sub-graphs, or trusses, within a network, thus, revealing the structural cohesion and tight-knit communities in complex graphs, facilitating the analysis of relationships and information flow in various fields, such as social networks, biology, and recommendation systems.
Keywords:
Mobile Agents Triangle Counting $k$-Truss Truss Decomposition Triangle Centrality Local Clustering Coefficient Time Complexity Memory Complexity Network Algorithms Distributed Algorithms
@@ 1 Introduction
Counting and listing triangles in a graph has received much attention in the last couple of decades as it serves as a building block of complex network analysis [54, 7]. The number of triangles in a graph is used for computing the clustering coefficient, one of the most used metrics for network analysis [54, 7], and triangle centrality [37, 1, 8]. Triangle counting also plays a pivotal role in the hierarchical decomposition of a graph such as truss decomposition [53] which is an important hierarchical subgraph structure in community detection [24, 2]. Triangle counting is used in solving many practical applications. Becchetti et al.  [6] used triangle counts in detecting web spam and estimating the content quality of a web page. Eckmann and Moses [18] have used the clustering coefficient in finding common topics on web pages. Other applications of triangle counting include query optimization in databases [5], link prediction in social network [51], and community detection in system biology [27].
@@
In this paper, we are interested in the triangle counting problem along with its applications in truss decomposition, computing triangle centrality and local clustering coefficient using autonomous agents (or robots) on anonymous graphs. Suppose we are given $n$ agents positioned initially on the nodes (each node has one agent) of an $n$-node anonymous graph $G$. The agents coordinate among themselves to solve the triangle counting problem such that each agent (at a node) outputs (i) node-based triangle count, (ii) edge-based triangle count and the total number of triangles in the graph (see the problem statements in Section 3).
@@
Our agent-based model has been gaining significant attention recently. For example, there have been some recent works on how to position the agents on nodes of the graph $G$ such that each agent’s position collectively form the maximal independent set (MIS) of $G$ [44, 42] or they identify a small dominating set [11] of $G$. Another related problem is of dispersion in which $k\leq n$ agents are positioned on $k$ different nodes of $G$, see [34] and the references therein. A solution to the dispersion problem guarantees that $k$ agents are positioned on $k$ different nodes; which is a requirement for the triangle counting problem defined in this paper. Exploration problem on graphs using mobile agents refers to solving a graph analytic task using one or more agents [16].
@@
In this work, we consider triangle counting in a simple, undirected, anonymous graph using mobile agents. The motivation stems from scenarios like private networks in the military or sensor networks in inaccessible terrain where direct access to the network is obstructed, but small battery-powered agents can navigate to learn network structures and their properties for overall network management. Prominent use of agents in network exploration can be seen in areas such as underwater navigation [15], network-centric warfare in military systems [36], modelling social network [56], studying social epidemiology [19] etc.
@@ 1.1 Our Contributions
We first enumerate the triangles in the graph $G$ and apply the triangle counting methodology to $3$ applications: (i) Truss Decomposition, (ii) Triangle Centrality, and (iii) Local Clustering Coefficient. Let $G$ be an $n$-node undirected and connected graph with the maximum degree $\Delta$ and diameter $D$. Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ ($c$ being an arbitrary constant) with the highest agent ID $\lambda\in[0,n^{c}]$, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration. Then, we solve the following problems.
@@
      Each agent $r_{i}$ can calculate and output the number of triangles adjacent to the vertex where $r_{i}$ is placed on, in $O(\Delta\log\lambda)$ rounds.
@@
      Each agent $r_{i}$ can calculate the number of triangles based on each of its adjacent edges in $O(\Delta\log\lambda)$ rounds.
@@
      Each agent $r_{i}$ can calculate the number of triangles in $G$ in $O(D\Delta\log\lambda)$ rounds.
@@
    Truss Decomposition - The Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta D\log\lambda)$ rounds with $O(\Delta\log n)$ bits of memory per agent.
  3. 3.
@@
1.2 Possible Applications
Triangle Counting and Truss Decomposition, in general, have various applications in community detection, social network analysis, biological networks, recommendation systems etc. Mobile agents on the other hand, equipped with truss decomposition capabilities, post dispersion, can dynamically identify and navigate through structurally significant regions of the graph. For example, an unmanned mobile agent can help prioritize areas within a large area that require prompt rescue and can carry rescue materials to difficult-to-access terrain. Similarly, robots carrying different agricultural materials (like fertilizers or pesticides) can dynamically arrange themselves and focus on specific areas that exhibit particular structural characteristics relevant to crop health. Truss decomposition can contribute to the autonomy of vehicles by helping them navigate efficiently through road networks or complex urban environments or helping them identify core areas in a city. In general, mobile robots can prioritize exploring areas with high truss density, indicating regions of potential interest or connectivity within the environment. In addition, its ability to modify itself in challenging scenarios improves fault tolerance and helps maintain the connectivity of the network.
@@ 2.1.1 Sequential Algorithms:
In 1985, Chiba and Nishizeki [13] proposed an algorithm for counting all triangles in a simple graph by computing the intersection of the neighborhoods of the adjacent vertices with time complexity $O(m^{\frac{3}{2}})$ where $m$ is the number of edges in the graph. Other sequential algorithms for triangle counting are based on vertex-iterator [47] and edge-iterator [26]. In vertex iterator based technique, it iterates over each vertex $v$ of the graph and intersects the adjacency list of each pair of the neighbors of $v$. In edge-iterator based technique, it iterates over each edge and intersects the adjacency list of its two endpoints.
@@ 2.2 Truss Decomposition
In [14], Cohen introduced and visualized truss as a relaxation to cliques in graphs and defined it to be a non-trivial, one-component sub-graph such that each edge is reinforced by at least $k-2$ pairs (for a $k$-truss) of edges making a triangle with that edge. He established the existence of polynomial time algorithms for identifying $k$-trusses within a given graph. Since then, trusses have become the subject of extensive investigation the problem has been studied in various models  [24, 12, 25, 29]. In general, truss decomposition is explored in two main contexts: a serialized version, primarily suitable for small to medium-sized graphs, and a parallel version designed for handling larger graphs. In [53], Yang et al. provided an improved serialized in-memory algorithm for computing $k-$truss in graphs of moderate sizes. In particular, their algorithm computed the $k-$trusses for all $k\geq 3$ in $O(m^{1.5})$ time using $O(m+n)$ memory space. They further proposed two I/O efficient algorithms for handling massive networks which are ill-equipped for the main memory of a single machine. To tackle the large memory requirement for massive graphs, several parallel implementations for truss decomposition were introduced. In [28], the authors implemented a parallel version of the serialized algorithm in [53] and used data structures adaptive to concurrent updates instead of hash tables. Sariyuce et al. in [46] employed the iterative h-index computation, formulated by [38], for nucleus decomposition and proved its convergence bounds. The truss decomposition problem is a particular case of the general nucleus decomposition problem. They developed parallel algorithms for both synchronous and asynchronous versions of the nucleus decomposition problem. In the synchronous version, the computations for specific iterations are made over a single snapshot of current values whereas in the asynchronous version, the latest values of each variable is used instantly in the current computation. Voegele et al. in [52] proposed a parallel graph-centric $k-$ truss decomposition and established the relation between a $k-$ truss and a $k-$core. When a $k-$ core computation proceeds $k-$ truss, a lot of edges from the graph may be waived off from being processed since a $k-$ truss is always contained in a $k-1$ core. Jian Wuet al. in [55], engineered both the serialized and parallel algorithms of  [53] and  [46] to reduce their memory usage by optimizing the underlying data structures and by using WebGraph. In [20], the authors implemented the truss decomposition on probabilistic graphs. They proposed an algorithm based on h-index updating and obtained an upper bound on the number of iterations for convergence. The efficiency of most of the algorithms is described with the aid of extensive experimental results.
@@
Agent-based Computations on Graphs: In [49], Sudo et al. considered the exploration problem with a single agent in undirected graphs. Starting from an arbitrary node, the agent has to explore all the nodes and edges in the graph and return to the starting node. The authors used a whiteboard model, that reduced the memory requirement per robot. In  [30], the authors explored the graph exploration problem based on Depth-First-Search and studied the trade-off between node-memory vs robot-memory. In [17], Dereniowski et al. proposed an algorithm for collective graph exploration with a team of $k$ ($k$ being polynomial size) agents in $O(D)$ time. They also obtained almost tight bounds on the asymptotic relation between exploration time and team size, for large $k$, in both the local and the global communication model. Further results on agent-based graph exploration and dispersion (spreading agents across the graph so that there are at-most one agent at each node) can be found in several recent papers [31, 32, 33, 39, 40, 43, 10].
@@ 3 Model and Problem Definitions
Mobile Agents (or Mobile Robots): We have a collection of $n$ agents =$\mathcal{R}=\{r_{1},r_{2},...,r_{n}\}$ residing on the nodes of the graph in such a way that each node is occupied by a distinct ID agent at the start (known as dispersed configuration in literature). Each agent has a unique ID in the range $[0,n^{c}]$ ($c$ is an arbitrary constant) and has $O(\Delta\cdot\log(n))$ bits to store information. An agent therefore cannot store the whole graph structure information within its limited memory. An agent retains its memory as long as needed and it can be updated as required. Two or more agents can be present (co-located) at a node or pass through an edge in $G$. However, an agent is not allowed to stay on an edge. An agent can recognise the port number through which it has entered and exited a node. The agents do not have any visibility beyond their (current) location at a node. An agent at node $v$ can only see the adjacent ports (connecting to edges) at $v$. Only the collocated agents at a node can sense each other and exchange information. An agent can transfer all the information stored in its memory in a single round. We interchangeably use the terms robots and agents.
@@
$\mathcal{R}$ Collection of $n$ mobile robots $\{r_{1},r_{2},\dots,r_{n}\}$
$\lambda$ ID of the robot with maximum ID
@@ 3.1 Problem Statements
Problem 1.
(Triangle Counting using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ agents, each of which initially placed distinctly at each node of $G$. We solve the following problems.
  1. 1.
@@
Problem 2.
(Truss Decomposition using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which are initially placed distinctly at each node of $G$. The $n$ autonomous agents coordinate among themselves to solve the Truss Decomposition Problem.
Problem 3.
(Triangle Centrality using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$. The $n$ autonomous agents coordinate among themselves to compute the Triangle Centrality $TC(v)$ for each node $v\in G$.
Problem 4.
(Local Clustering Coefficient using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$. The $n$ autonomous agents coordinate among themselves to compute the Local Clustering Coefficient, $LLC(v)$ for each node $v\in G$.

### v2

@@ Abstract
Triangle counting in a graph is a fundamental problem with wide-ranging applications across various domains. It is crucial for understanding graph structure and serves as a basis for more advanced graph analytics. One key application of triangle counting is truss decomposition, a technique that helps in identifying maximal, highly interconnected sub-graphs, or trusses, within a network, revealing the structural cohesion and tight-knit communities in complex graphs, facilitating the analysis of relationships and information flow in various fields, such as social networks, biology, and recommendation systems. Utilizing mobile agents or robots for tasks like truss decomposition and clustering coefficient calculations can be particularly advantageous in decentralized environments with limited, unreliable, or compromised communication systems. In such scenarios, these agents can carry out local computations, bypassing the need for extensive communication channels. This approach proves valuable in contexts such as disaster response, urban management, and military operations, where broadcast communication may be impractical.
In this paper, we first address the triangle counting problem in an arbitrary anonymous graph using mobile agents. This method is then extended as a subroutine to solve the truss decomposition problem and compute the triangle centrality metric and local clustering coefficient for each node. Our approach utilizes $n$ autonomous mobile agents, each initially positioned at a different node of the $n$ node graph. These agents coordinate to collaboratively first solve the triangle enumeration problem and then truss decomposition and compute the triangle centrality and local clustering coefficient metrics. In line with existing research, we assume a synchronous system where all agents execute their tasks concurrently, allowing time complexity to be measured in rounds. The graph is anonymous, where the nodes lack any labels or IDs. However, the agents possess distinct IDs and some limited memory. Although the agents can perform extensive local computations, they can only communicate when they occupy the same node. Our primary goal is to design algorithms that minimize both the time required (which also reflects the number of agent movements) for triangle counting and the memory usage of each agent. Furthermore, we demonstrate how the triangle counts obtained through the mobile agent approach can be leveraged to solve the truss decomposition, triangle centrality, and local clustering coefficient problems.
Keywords Agent-Based Systems, Autonomous Agents, Distributed Robot Systems, Multi-Robot Systems, Mobile Agents, Triangle Counting, $k$-Truss, Truss Decomposition, Triangle Centrality, Local Clustering Coefficient, Time Complexity, Network Algorithms, Distributed Algorithms
@@ 1 Introduction
Counting and listing triangles in a graph has received much attention in the last couple of decades as it serves as a building block of complex network analysis [2, 3]. The triangle count in a graph is used for computing clustering coefficient, one of the most used metrics for network analysis [2, 3], and triangle centrality [4, 5, 6]. Triangle counting also plays a pivotal role in the hierarchical decomposition of a graph such as truss decomposition [7] which is an important hierarchical subgraph structure in community detection [8, 9]. Triangle counting is useful in many practical applications. Becchetti et al.  [10] used triangle counts in detecting web spam and estimating the content quality of a web page. Eckmann and Moses [11] have used the clustering coefficient in finding common topics on web pages. Other applications of triangle counting include query optimization in databases [12], link prediction in social network [13], and community detection in system biology [14].
@@
In this paper, we are interested in the triangle counting problem along with its applications in truss decomposition, computing triangle centrality and local clustering coefficient using autonomous agents on anonymous graphs. Suppose we are given $n$ agents positioned initially on the nodes (each node has one agent) of an $n$-node anonymous graph $G$. The agents coordinate among themselves to solve the triangle counting problem such that each agent (at a node) outputs (i) a node-based triangle count, (ii) an edge-based triangle count, and the total number of triangles in the graph (see the problem statements in Section 3).
@@
Our agent-based model has been gaining significant attention recently. For example, there have been some recent work on how to position the agents on nodes of the graph $G$ such that each agent’s position collectively forms the maximal independent set (MIS) of $G$ [15, 16] or they identify a small dominating set [17] of $G$. Recently, a similar mobile agent model has been used to explore the complexity of constructing a breadth-first search tree [18] and a minimum spanning tree in arbitrary anonymous graphs [19]. Another related problem is of dispersion in which $k\leq n$ agents have to position themselves at $k$ different nodes of $G$, see [20] and the references therein. A solution to the dispersion problem guarantees that $k$ agents are positioned on $k$ different nodes; this is a requirement for the triangle counting problem defined in this paper. Exploration problems on graphs using mobile agents refer to solving a graph analytic task using one or more agents [21].
@@
In this work, we consider triangle counting in a simple, undirected, anonymous graph using mobile agents. The motivation comes from scenarios such as private networks in the military or sensor networks in inaccessible terrain where direct access to the network is obstructed, but small battery-powered agents can navigate to learn network structures and their properties for overall network management. The prominent use of agents in network exploration can be seen in areas such as underwater navigation [22], network-centric warfare in military systems [23], modeling social networks [24], studying social epidemiology [25] etc. In addition, triangle counting can be effectively applied to drone-based Social Network Analysis (SNA) in scenarios such as drone-assisted event monitoring at public gatherings, protests, or concerts. For example, as demonstrated in [26], drones can identify individuals in close proximity, such as those not wearing masks, and issue social distancing alerts through speakers. In addition, drones can collect data on interactions between individuals and triangle counting can be used to identify closely knit social groups, discover group dynamics, detect leaders, or monitor potentially suspicious clusters in real time, enhancing situational awareness and crowd management.
Furthermore, mobile agents, such as drones, can play a crucial role in Search and Rescue (SAR) operations [27, 28, 29, 30]. In SAR missions, multiple drones can cover large areas, locate missing persons, and maintain communication with the command centre. Drones can form resilient triangular communication networks where triangle counting helps identify strong communication links. This ensures robust communication paths between drones and the base station, facilitating quick information sharing.
Drones are also being used for Wireless Sensor Network (WSN) deployment in remote or difficult-to-access areas, especially for environmental monitoring [31, 32, 33, 34]. By identifying dense clusters of sensors or drones, triangle counting enhances network robustness. It helps detect areas of high data redundancy, optimize sensor placement, and improve communication resilience.
@@ 1.1 Our Contributions
We first enumerate the triangles in the graph $G$ and apply the triangle counting methodology to three applications: (i) Truss Decomposition, (ii) Triangle Centrality, and (iii) Local Clustering Coefficient. Let $G$ be an $n$-node simple, undirected, anonymous and connected graph with the maximum degree $\Delta$ and diameter $D$. Let $n$ mobile agents with distinct IDs with the highest agent ID $\lambda$, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration. Then, we solve the following problems.
@@
      Each agent $r_{i}$ can calculate and output the number of triangles containing the vertex where $r_{i}$ is placed on, in $O(\Delta\log\lambda)$ rounds.
@@
      Each agent $r_{i}$ can calculate the number of triangles based on each edge of $G$ in $O(\Delta\log\lambda)$ rounds.
@@
      Each agent $r_{i}$ can calculate the total number of triangles in $G$ in $O(D\Delta\log\lambda)$ rounds.
@@
    Applications
    • $\bullet$
      Truss Decomposition - The Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta\log\lambda+mD)$ rounds.
    • $\bullet$
@@
    The memory requirement for each agent is $O(\Delta\log n)$ bits.
@@ 2.1.1 Sequential Algorithms:
In 1985, Chiba and Nishizeki [35] proposed an algorithm for counting all triangles in a simple graph by computing the intersection of the neighbourhoods of the adjacent vertices with time complexity $O(m^{\frac{3}{2}})$ where $m$ is the number of edges in the graph. Other sequential algorithms for triangle counting are based on vertex-iterator [36] and edge-iterator [37]. In the vertex iterator based technique, it iterates over each vertex $v$ of the graph and intersects the adjacency list of each pair of the neighbours of $v$. In the edge-iterator based technique, it iterates over each edge and intersects the adjacency list of its two endpoints.
@@ 2.2 Truss Decomposition
In [48], Cohen introduced and visualized truss as a relaxation to cliques in graphs and defined it to be a non-trivial, one-component sub-graph such that each edge is reinforced by at least $k-2$ pairs (for a $k$-truss) of edges, making a triangle with that edge. He established the existence of polynomial time algorithms for identifying $k$-trusses within a given graph. Since then, trusses have become the subject of extensive investigation; the problem has been studied in various models  [8, 49, 50, 51]. In general, truss decomposition is explored in two main contexts: a serialized version, primarily suitable for small to medium-sized graphs, and a parallel version designed for handling larger graphs. In [7], Yang et al. provided an improved serialized in-memory algorithm for computing $k-$truss in graphs of moderate sizes. In particular, their algorithm computed the $k-$trusses for all $k\geq 3$ in $O(m^{1.5})$ time using $O(m+n)$ memory space. They further proposed two I/O efficient algorithms for handling massive networks, which are ill-equipped for the main memory of a single machine. To tackle the large memory requirement for massive graphs, several parallel implementations for truss decomposition were introduced. In [52], the authors implemented a parallel version of the serialized algorithm in [7] and used data structures adaptive to concurrent updates instead of hash tables. Sariyuce et al. in [53] employed the iterative h-index computation, formulated by [54], for nucleus decomposition and proved its convergence bounds. The truss decomposition problem is a particular case of the general nucleus decomposition problem. They developed parallel algorithms for both synchronous and asynchronous versions of the nucleus decomposition problem. In the synchronous version, the computations for specific iterations are made over a single snapshot of current values, whereas in the asynchronous version, the latest values of each variable is used instantly in the current computation. Voegele et al. in [43] proposed a parallel graph-centric $k-$ truss decomposition and established the relation between a $k-$ truss and a $k-$core. When a $k-$ core computation proceeds $k-$ truss, a lot of edges from the graph may be waived off from being processed since a $k-$ truss is always contained in a $k-1$ core. Jian Wuet al. in [55], engineered both the serialized and parallel algorithms of  [7] and  [53] to reduce their memory usage by optimizing the underlying data structures and by using WebGraph. In [56], the authors implemented the truss decomposition on probabilistic graphs. They proposed an algorithm based on h-index updating and obtained an upper bound on the number of iterations for convergence. The efficiency of most of the algorithms is described with the aid of extensive experimental results.
2.3 Agent-based Computations on Graphs:
@@
In [57], Sudo et al. considered the exploration problem with a single agent in undirected graphs. Starting from an arbitrary node, the agent has to explore all the nodes and edges in the graph and return to the starting node. The authors used a whiteboard model, that reduced the memory requirement per agent. In  [58], the authors explored the graph exploration problem based on Depth-First-Search and studied the trade-off between node-memory vs. agent-memory. In [59], Dereniowski et al. proposed an algorithm for collective graph exploration with a team of $k$ ($k$ being polynomial size) agents in $O(D)$ time. They also obtained almost tight bounds on the asymptotic relation between exploration time and team size, for large $k$, in both the local and the global communication model. Further results on agent-based graph exploration and dispersion (spreading agents across the graph so that there is at-most one agent at each node) can be found in several recent papers [60, 61, 62, 63, 64, 65, 66].
@@ 3 Model and Problem Definitions
Mobile Agents: We have a collection of $n$ agents =$\mathcal{R}=\{r_{1},r_{2},...,r_{n}\}$ residing on the nodes of the graph in such a way that each node is occupied by a distinct ID agent at the start (known as dispersed configuration in literature). Each agent has a unique ID (we bound the IDs of the agents by $n^{c}$, for some large constant $c$, to accommodate the ID space of the agents within an $O(\log n)$ bit ID field) and has a limited memory to store information. An agent cannot store the whole graph structure information within its memory. An agent retains its memory as long as needed, and it can be updated as required. Two or more agents can be present (co-located) at a node or pass through an edge in $G$. However, an agent is not allowed to stay on edge. An agent can recognise the port number through which it has entered and exited a node. The agents do not have any visibility beyond their (current) location at a node. An agent at node $v$ can only see the adjacent ports (connecting to edges) at $v$. Only the collocated agents at a node can sense each other and exchange information. Every agent has prior knowledge of the highest degree of the graph $\Delta$ and its diameter $D$. The agents are also aware of the highest ID among themselves, denoted as $\lambda$. An agent can transfer all the information stored in its memory in a single round.
@@
$\mathcal{R}$ Collection of $n$ mobile agents $\{r_{1},r_{2},\dots,r_{n}\}$
$\lambda$ ID of the agent with maximum ID
@@ 3.1 Problem Statements
Triangle Counting using Mobile Agents: Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$. We solve the following problems.
  1. (a)
@@
Truss Decomposition, Triangle Centrality and Local Clustering Coefficient: Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$. The $n$ autonomous agents coordinate among themselves to solve the (i) Truss Decomposition Problem and compute (ii) Triangle Centrality, and (iii) Local Clustering Coefficient of a given node.